OpenStudy (anonymous):

The height of a parallelogram is 8 units more than the base. The measure of the base of the parallelogram is 12 units. What is the area, in square units, of the parallelogram? Answer 48 96 120 240

OpenStudy (anonymous):

its 240

OpenStudy (anonymous):

it was 240

OpenStudy (anonymous):

thank-you guys if you don't mind how did you work that out i think i worked it out wrong...

OpenStudy (anonymous):

i did this question for my unit test, its right; H = 8 + B B = 20 H = 8 + 20 Area of a parallelogram = BH A = 12(20) A = 240

OpenStudy (anonymous):

i mean B = 12 H = 8 + 12 H = 20 A = 12(20) A = 240

OpenStudy (anonymous):

got it! ok i just had the formula wrong mmk thanks!

OpenStudy (anonymous):

i know what you meant haha thanks again =)